OpenStudy (mcoban):

An electron has an uncertainty in its position of 552pm. What is the uncertainty in its velocity?

OpenStudy (jfraser):

plug into the uncertainty equation: \[\Delta x*\Delta mv \ge \frac{h}{4\pi}\] rearrange and solve for Dv